OpenStudy (kj4uts):

1. A flask contains 4.00 g of solid sodium hydroxide. Some water is added to the flask to dissolve the sodium hydroxide, and enough water is added to fill the flask to a volume of 50.0 mL. What volume of 1.00 M HCl is needed to neutralize the sodium hydroxide? A. 50.0 mL B. 100.0 mL C. 200.0 mL 2. A beaker contains 100.0 mL of a 0.100 M solution of sulfuric acid. When 100.0 mL water is added to this solution, what is the concentration of the diluted sulfuric acid? A. 0.0250 M B. 0.0500 M C. 0.200 M 3. How many moles of NaOH are required to react completely with 1.00 mole of nitric acid? A. 1.00 mol B. 0.500 mol C. 2.000 mol Please explain. Thank you!

NvidiaIntely (nvidiaintely):

Sorry I do not know how to explain this. However I found that @Kevin, is great at this stuff. He will be online soon, so stay online. yw

NvidiaIntely (nvidiaintely):

@Kevin, make sure you explain it to this guy.

OpenStudy (sapphiremoon):

I could do this, but it's pretty late in my part of the world and with my schedule, so I can get on this first thing in the morning if no one else answers, seem ok @KJ4UTS?

OpenStudy (kj4uts):

@SapphireMoon no problem, I appreciate any help you can give me. This is a new unit for me and im having some difficulty understanding it. Thanks again :)

OpenStudy (jiteshmeghwal9):

Answer 1 :- concentration of NaOH after dissolution of 50 ml of water in it = \[\text{Molarity }= \frac{\text{no. of moles of solute} }{\text{volume of solution in liters}}\] number of moles of NaOH= \(\Large\frac{4}{40}\) volume of solution in liters = \(\Large \frac{50}{1000}\) L concentration of NaOH = \(\Large \frac{4/40}{50/1000}\) M Remember the fact that during Neutralization reaction Milli-equivalents of acid = Milliequivalents of base (number of milliequivalents = Molarity * n-factor * volume of solution in milliliters)

OpenStudy (jiteshmeghwal9):

n-factor for HCl =1 n-factor for NaOH =1 Let the volume of HCl required for neutralization be 'V' ml Milliequivalents of HCl = Milliequivalents of NaOH 1M * 1 * V= \(\Large \frac{4/40}{50/1000}\) * 1 * 50 .......(1) solve equation (1) yo get 'V'. remember u will gt volume in milliliters

OpenStudy (jiteshmeghwal9):

Answer :- 2 During dilution number of moles of the solute do not alter so \[M_1 V_1=M_2 V_2\] \(M_1\) = initial concentration \(M_2\)=final concentration \(V_1\)=Initial volume of solution \(V_2\)=final volume of solution can u apply this rule now to find answer for question 2 ?

OpenStudy (jiteshmeghwal9):

Answer :- As i told u earlier that during neutralization reaction (Milli-equivalents of acid = Milliequivalents of base) (milliequivalent of a compound = no. of moles of that substance * n-factor) milliequivalents of nitric acid = 1*1=1 (as both no. of moles & n-factor of \(HNO_3\) are 1) Let the number of moles of NaOH required for neutralization of \(HNO_3\) be 'n'. milliequivalents of NaOH = n*1 (as no. of moles of NaOH is n & n-factor is 1) Therefore, milliequivalents of nitric acid=milliequivalents of NaOH 1*1=n*1

OpenStudy (kevin):

1) M = gr/Mr x 1000/p M = 4/40 x 1000/50 M = 2 M1 x V1 = M2 x V2 2 x 50 = 1 x V2 V2 = 100 2) M1 x V2 = M2 x V2 (just like the question on number 1) 3) HNO3 + NaOH -> NaNO3 + H2O HNO3 -> H^+ + NO3^- -> valency of H+ = 1 NaOH -> Na^+ + OH^- -> valency of OH^- = 1 M1 x V1 x valency = M2 x V2 x valency mole1 x valency = mole2 x valency 1 x 1 = mole2 x 1